# U-shaped and Inverted-U Scaling behind Emergent Abilities of Large Language Models

### v1

@@ Appendix D Scaling Trend by Question Difficulty Level for Non-emergent Tasks
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure A18: The accuracy, binary Brier Score, and scaling trend on the abstract narrative understanding dataset in BIG-bench (Srivastava et al., 2023).
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure A19: The accuracy, binary Brier Score, and scaling trend on the ARC dataset (Clark et al., 2018).
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure A20: The accuracy, binary Brier Score, and scaling trend on the HellaSwag dataset (Zellers et al., 2019).
Appendix E Scaling Trend by Question Difficulty Level on Accuracy
@@
We apply the same procedure as in Sec. 2 with accuracy as the performance measure instead of binary Brier Score. Specifically, we calculate question difficulty level using average accuracy over models before the emergence threshold and plot the scaling trend on accuracy for each difficulty group, as shown in Fig. A21. We still observe U-shaped scaling for the hard question groups and inverted-U scaling for the easy question groups followed by steady improvement after the emergence threshold. However, the scaling trend on accuracy is much less pronounced than that on binary Brier Score. Accuracy on the arithmetic and Persian-QA datasets exhibits nearly flat scaling up to the emergence threshold. We attribute this to the crude nature of accuracy as a performance measure.
Appendix F More Discussions on Slice-and-Sandwich
F.1 Robustness Analysis
We present the robustness analysis of Slice-and-Sandwich regarding (1) the choice of order and (2) lower model size cutoff for the training set.
F.1.1 Effect of Polynomial Degree
@@
Fig. A22 shows the polynomial fit of binary Brier Score for the easy group with degree=3, 5, and 7, and for the hard group with degree=2, 4, and 6. Note that we consider only polynomials of odd and even degrees for the hard and easy question groups, respectively. This prior knowledge reflects the observation that performance of the easy question group initially improves with scale, the performance of the hard question group initially declines with scale, whereas the performance of both groups increases with scale past the emergence threshold. In general, there is a bias-variance tradeoff: a polynomial fit of a higher degree has higher recall but lower precision. The polynomial fit of a higher degree might be over-sensitive to noises in the training data, while the polynomial fit of a lower degree might lack the flexibility to capture the turning points in data.
@@
F.1.2 Effect of Model-Size Threshold for Train-Test Split
@@
The forecast is reasonably robust to the train-test split. All capture the trend and display a similar shape to our original choice of train-test split threshold except for the case where threshold$=1.3$ and degree$=5$ for the MMLU dataset. The polynomial fit of degree 3 for Persian QA is too flat compared to data for all three thresholds and gets flatter as the threshold goes down. We leave it to future work to provide better guidelines as to the least upper bound of training data model size that still allows us to confidently predict the onset of emergent abilities.
F.2 Hard Lift - A Simple Alternative Pipeline
@@
As an alternative to Slice-and-Sandwich, we provide an even simpler pipeline called Hard-Lift. Specifically, we take the polynomial fit of degree 2 on binary Brier Score for the hard question group from Slice-and-Sandwich and lift it by a constant so the fitted binary Brier Score at the training set model size upper bound is equal to the true average. We use this to forecast the binary Brier Score of models past the emergence threshold. We then transform this predicted binary Brier Score back to predicted accuracy via the $G(\cdot)$ function as in Slice-and-Sandwich.
Fig. A26 shows the results of Hard-Lift under different model-size thresholds for train-test split as in Sec. F.1.2. Hard-Lift performs better than the baseline for MMLU (Fig. 26(a)) and Persian-QA (Fig. 26(c)) datasets, but worse than baseline for the arithmetic dataset. We believe this result reinforces our claim that analyzing difficulty-stratified scaling trends enables more explainable prediction of emergent abilities.
Appendix G Preliminary Analysis for String-Match Tasks
(a) Exact match accuracy vs. effective model size (M).
(b) Token edit distance vs. effective model size (M).
(c) token edit distance vs. effective model size (M) with group number $G=3$.
(d) Modified cosine similarity vs. effective model size (M) with group number $G=3$.
Figure A27: The exact match accuracy, token edit distance (TED), and cosine similarity score vs. effective model size (M) on the word unscramble dataset in BIG-bench (Srivastava et al., 2023).
@@ Appendix G Preliminary Analysis for String-Match Tasks
We argue that TED is not a good measure of progress on a string match task. (1) It does not differentiate between easy and hard questions well. Fig.  27(c) shows that performance measured by TED on all three question groups is close for all model sizes above $0.5$. A harder group’s TED may be higher or lower than an easier group’s. (2) It is not very correlated with exact match, the traditional metric that people are probably ultimately interested in ( 28(a)).
@@
Interestingly, performance measured by MCS for all question groups exhibits inverted-U scaling followed by steady improvement. The only differences are the model size at which scaling reverts from inverse scaling to standard scaling and also how fast performance goes up/down. However, Fig. 28(b) shows that MCS is also poorly correlated with the exact match. Even if we can precisely predict the MCS of models above the emergence threshold, conversion back to exact match accuracy will be too noisy to be useful. We hope this section illustrates potential avenues for future work.
Appendix H Broader Impact
H.1 Potential Positive Impacts
@@
H.2 Potential Negative Impacts

### v2

@@ Appendix D Scaling Trend by Question Difficulty Level for Non-emergent Tasks
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure A18: The accuracy, TC Brier Score, and scaling trend on the abstract narrative understanding dataset in BIG-bench (Srivastava et al., 2023).
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure A19: The accuracy, TC Brier Score, and scaling trend on the ARC dataset (Clark et al., 2018).
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure A20: The accuracy, TC Brier Score, and scaling trend on the HellaSwag dataset (Zellers et al., 2019).
Appendix E More MMLU Sample Questions Exhibiting U-Shaped and Inverted-U Scaling
Table A3: Six questions in the easiest question group of the MMLU, with the correct choices underlined and each question’s TC Brier Scores by Pythia-70m ($M=-0.90$), Pythia-160m ($M=-0.54$), and Pythia-1b ($M=0.40$) displayed. Pythia of all sizes are pre-trained on the same data and in the same order (Biderman et al., 2023), so the model parameter number is the only variable that affects log compute. The six questions demonstrate inverted-U scaling: Pythia-160m performs better than Pythia-70m, while Pythia-1b performs worse than Pythia-160m.
Description Choices Pythia -70m Pythia -160m Pythia -1b
(high school mathematics, id = 142) Simplify the following expression: $(9x^{9}+7x^{8}+4x^{7})+(x^{11}+x^{9}+2x^{7}+3x^{3}+5x+8).$ Express your answer as a polynomial with the degrees of the terms in decreasing order. A. $x^{11}+2x^{9}+2x^{8}$ B. $x^{11}-6x^{8}+6x^{7}+3x^{3}+5x+8$ C. $x^{1}1+10x^{9}+7x^{8}+6x^{7}3x^{3}$ $+5x+8$ D. $x^{11}+10x^{9}+7x^{8}+6x^{7}+3x^{3}$ $+5x+8$ A. 0.07 B. 0.20 C. 0.53 D. 0.20 A. 0.06 B. 0.06 C. 0.16 D. 0.72 A. 0.18 B. 0.28 C. 0.28 D. 0.26
(anatomy, id = 120) Which of the following structures is part of the small intestine? A. Ascending colon B. Cecum C. Ileum D. Sigmoid colon A. 0.04 B. 0.32 C. 0.32 D. 0.32 A. 0.16 B. 0.26 C. 0.43 D. 0.16 A. 0.18 B. 0.26 C. 0.27 D. 0.29
(high school physics, id = 124) Which of the following conditions are necessary for an object to be in static equilibrium? I. The vector sum of all torques on the object must equal zero. II. The vector sum of all forces on the object must equal zero. III. The sum of the object’s potential and kinetic energies must be zero. A: I only B. II only C. III only D. I and II only A. 0.08 B. 0.61 C. 0.08 D. 0.22 A. 0.11 B. 0.19 C. 0.19 D. 0.51 A. 0.15 B. 0.23 C. 0.28 D. 0.34
(high school macroeconomics, id = 104) In order to reduce or eliminate crowding out expansionary fiscal policy can be accompanied by A: an increase in government spending B. a decrease in investment C. expansionary monetary policy D. contractionary monetary policy A. 0.06 B. 0.06 C. 0.44 D. 0.44 A. 0.14 B. 0.09 C. 0.63 D. 0.14 A. 0.29 B. 0.32 C. 0.24 D. 0.16
(miscellaneous, id = 559) Which of these is made from cacao seeds? A. marzipan B. soy sauce C. chocolate D. anchovies A. 0.37 B. 0.13 C. 0.37 D. 0.13 A. 0.16 B. 0.10 C. 0.71 D. 0.04 A. 0.15 B. 0.32 C. 0.32 D. 0.22
(professional accounting, id = 65) A project should be accepted if the present value of cash flows from the project is A. Equal to the initial investment. B. Less than the initial investment. C. Greater than the initial investment. D. Equal to zero. A. 0.37 B. 0.37 C. 0.13 D. 0.13 A. 0.26 B. 0.26 C. 0.43 D. 0.06 A. 0.22 B. 0.29 C. 0.34 D. 0.15
Table A4: Six questions in the hardest question group of the MMLU, with the correct choices underlined and each question’s TC Brier Scores by Pythia-70m ($M=-0.90$), Pythia-160m ($M=-0.54$), and Pythia-1b ($M=0.40$) displayed. The six questions demonstrate U-shaped scaling: Pythia-160m performs worse than Pythia-70m, while Pythia-1b performs better than Pythia-160m.
Description Choices Pythia -70m Pythia -160m Pythia -1b
(moral scenarios, id = 220) For which of these two scenarios does the main character (who uses I/me/my) do something clearly morally wrong, according to ordinary moral standards in the US as of 2020? Scenario 1 | As my parents slept, I played Tic-tac-toe. Scenario 2 | I slid down the slide with no one on it. A. Wrong, Wrong B. Wrong, Not wrong C. Not wrong, Wrong D. Not wrong, Not wrong A. 0.17 B. 0.17 C. 0.17 D. 0.48 A. 0.17 B. 0.29 C. 0.47 D. 0.06 A. 0.18 B. 0.22 C. 0.31 D. 0.29
(high school mathematics, id = 159) At Academic Academy, to pass an algebra test you must score at least $80\\ \%$. If there are 35 problems on the test, what is the greatest number you can miss and still pass? A. 7 B. 28 C. 35 D. 8 A. 0.17 B. 0.17 C. 0.17 D. 0.48 A. 0.04 B. 0.07 C. 0.33 D. 0.55 A. 0.08 B. 0.34 C. 0.33 D. 0.24
(college biology, id = 44) Natural enemies have been implicated as a strong selective force for all of the following EXCEPT A. aposematic coloration B. chemical defenses C. masting (synchronous fruiting) D. lekking behavior A. 0.17 B. 0.48 C. 0.17 D. 0.17 A. 0.18 B. 0.49 C. 0.29 D. 0.04 A. 0.37 B. 0.26 C. 0.21 D. 0.16
(conceptual physics, id = 219) Light reflecting from a smooth surface undergoes a change in A. frequency. B. wavelength. C. All of these. D. None of these. A. 0.30 B. 0.30 C. 0.30 D. 0.11 A. 0.57 B. 0.21 C. 0.21 D. 0.02 A. 0.21 B. 0.34 C. 0.27 D. 0.18
(astronomy, id = 18) Which of the following is/are common feature(s) of all fresh (i.e. not eroded) impact craters formed on solid surfaces: A. ejecta B. raised rims C. central peaks D. A and B only A. 0.48 B. 0.17 C. 0.17 D. 0.17 A. 0.82 B. 0.11 C. 0.04 D. 0.02 A. 0.36 B. 0.31 C. 0.19 D. 0.14
(professional accounting, id = 70) Grant Co.’s sales budget shows the following projections for the year ending December 31: Quarter Units First 30000 Second 40000 Third 22500 Fourth 27500 Total 120000 Inventory at the beginning of the year was budgeted at 9000 units. The quantity of finished goods inventory at the end of each quarter is to equal 30% of the next quarter’s budgeted sales of units. What amount should the production budget show for units to be produced during the first quarter? A. 36000 B. 33000 C. 24000 D. 12000 A. 0.44 B. 0.44 C. 0.06 D. 0.06 A. 0.17 B. 0.28 C. 0.46 D. 0.10 A. 0.24 B. 0.25 C. 0.32 D. 0.19
This section provides more qualitative results of MMLU questions to further illustrate the U-shaped and inverted-U scaling. Table A3 presents six questions in the easiest question group (difficulty level 1) and their performance on four LLMs measured by TC Brier Score. In particular, we adopt Pythia-70m, Pythia-160m, Pythia-1b, and Qwen1.5-14b, ordered from small to large log compute. All Pythia models are pre-trained on the same corpus (Biderman et al., 2023). As a result, only the model parameter size will affect Pythia’s log compute. As shown in Table A3, these questions demonstrate an apparent inverted-U scaling: Pythia-160m obtains a higher TC Brier Score compared with Pythia-70m by assigning a relatively higher confidence to the target choice (class), while Pythia-1b underperforms Pythia-160m. Table A4 presents six questions in the hardest question group (difficulty level 10) of MMLU and their performances of three Pythia models and Qwen1.5-14b, which exhibit the U-shaped scaling: Pythia-160m gets worse than Pythia-70m by assigning low relative confidence to the target choice, while Pythia-1b performs better than Pythia-160m. Notably, Pythia-1b’s performance can vary relative to Pythia-70m, depending on the question.
Appendix F Scaling Trend by Question Difficulty Level on Accuracy
@@
We apply the same procedure as in Sec. 2 with accuracy as the performance measure instead of the TC Brier Score. Specifically, we calculate question difficulty level using average accuracy over models before the emergence threshold and plot the scaling trend on accuracy for each difficulty group, as shown in Fig. A21. We still observe clear inverted-U scaling for the easiest question group followed by steady improvement after the emergence threshold. However, hard question groups no longer exhibit a clear U-shaped scaling. Specifically, accuracy performance of hard question groups tends to stagnate after the initial performance drop. For instance, all three datasets’ hardest hard question groups stagnate at near-zero accuracy, lower than the random guess. The worse-than-random performance can be explained by distracting questions, as discussed in Sec. 3.2. On the other hand, the mitigated U-shaped scaling might be due to the fact that accuracy does not capture the change in the model’s confidence level in the target class. In other words, the accuracy-based procedure cannot capture the models’ learning process of first being distracted by questions and gradually overcoming the distraction, because the models’ accuracies are all around zero.
Appendix G More Discussions on Slice-and-Sandwich
G.1 Robustness Analysis
We present the robustness analysis of Slice-and-Sandwich regarding (1) the choice of order, (2) lower model log compute cutoff for the training set, and (3) group number $G$.
G.1.1 Effect of Polynomial Degree
@@
Fig. A22 shows the polynomial fit of TC Brier Score for the easy group with degree=3, 5, and 7, and for the hard group with degree=2, 4, and 6. Note that we consider only polynomials of odd and even degrees for the hard and easy question groups, respectively. This prior knowledge reflects the observation that performance of the easy question group initially improves with scale, the performance of the hard question group initially declines with scale, whereas the performance of both groups increases with scale past the emergence threshold. In general, there is a bias-variance tradeoff: a polynomial fit of a higher degree has higher recall but lower precision. The polynomial fit of a higher degree might be over-sensitive to noises in the training data, while the polynomial fit of a lower degree might lack the flexibility to capture the turning points in data.
@@
G.1.2 Effect of Log Compute Threshold for Train-Test Split
@@
The forecast is reasonably robust to the train-test split. All capture the trend and display a similar shape to our original choice of train-test split threshold except for the case where threshold$=1.3$ and degree$=5$ for the MMLU dataset. The polynomial fit of degree 3 for Persian QA is too flat compared to data for all three thresholds and gets flatter as the threshold goes down. We leave it to future work to provide better guidelines as to the least upper bound of training data model log compute that still allows us to confidently predict the onset of emergent abilities.
G.1.3 Effect of Group Number
(a) MMLU with group number $G$ being $3$, $5$, and $7$.
(b) Arithmetic with group number $G$ being $3$, $5$, and $7$.
(c) Persian-QA with group number $G$ being $3$, $5$, and $7$.
Figure A26: Slice-and-Sandwich’s results of accuracy-based scaling law under different group numbers $G$.
Fig. A26 shows the fitted scaling trends when splitting questions into different numbers of groups. We show group number $G=3,5,$ and $7$. Following the same procedure and degree parameter in the main paper, easier question groups, such as the groups of difficulty level 1 to 3 for $G=7$, are fitted by polynomial regression of degree=5 due to observed inverted-U scaling; harder question groups are fitted by degree=2 due to U-shaped scaling. Then Eq. 5 is modified to take the average of Brier-based fitting trends of all but the medium group and project the acquired Brier-based scaling law to the accuracy-based one. Slice-and-Sandwich shows its robustness under $G$. The robustness comes from similar fitting results among the same scaling types, such as inverted-U scaling, resulting in a similar final scaling law after taking their averages.
G.2 Hard Lift - A Simple Alternative Pipeline
@@
As an alternative to Slice-and-Sandwich, we provide an even simpler pipeline called Hard-Lift. Specifically, we take the polynomial fit of degree 2 on TC Brier Score for the hard question group from Slice-and-Sandwich and lift it by a constant so the fitted TC Brier Score at the training set model log compute upper bound is equal to the true average. We use this to forecast the TC Brier Score of models past the emergence threshold. We then transform this predicted TC Brier Score back to predicted accuracy via the $G(\cdot)$ function as in Slice-and-Sandwich.
Fig. A28 shows the results of Hard-Lift under different log compute thresholds for train-test split as in Sec. G.1.2. Hard-Lift performs better than the baseline for MMLU (Fig. 28(a)) and Persian-QA (Fig. 28(c)) datasets, but worse than baseline for the arithmetic dataset. We believe this result reinforces our claim that analyzing difficulty-stratified scaling trends enables more explainable prediction of emergent abilities.
Appendix H Preliminary Analysis for String-Match Tasks
(a) Exact match accuracy vs. log compute (M).
(b) Token edit distance vs. log compute (M).
(c) token edit distance vs. log compute (M) with group number $G=3$.
(d) Modified cosine similarity vs. log compute (M) with group number $G=3$.
Figure A29: The exact match accuracy, token edit distance (TED), and cosine similarity score vs. log compute (M) on the word unscramble dataset in BIG-bench (Srivastava et al., 2023).
@@ Appendix H Preliminary Analysis for String-Match Tasks
We argue that TED is not a good measure of progress on a string match task. (1) It does not differentiate between easy and hard questions well. Fig.  29(c) shows that performance measured by TED on all three question groups is close for all model log computes above $0.5$. A harder group’s TED may be higher or lower than an easier group’s. (2) It is not very correlated with exact match, the traditional metric that people are probably ultimately interested in ( 30(a)).
@@
Interestingly, performance measured by MCS for all question groups exhibits inverted-U scaling followed by steady improvement. The only differences are the model log compute at which scaling reverts from inverse scaling to standard scaling and also how fast performance goes up/down. However, Fig. 30(b) shows that MCS is also poorly correlated with the exact match. Even if we can precisely predict the MCS of models above the emergence threshold, conversion back to exact match accuracy will be too noisy to be useful. We hope this section illustrates potential avenues for future work.
Appendix I Broader Impact
I.1 Potential Positive Impacts
@@
I.2 Potential Negative Impacts